OpenStudy (anonymous):

Jessi rents a moving van for $75 and must pay $2 per mile. The following week, Alex rents the same van, is charged $80 for the rental and $1.50 per mile. If they each paid the same amount and drove the same number of miles, how far did they each travel?

OpenStudy (skyz):

Uh shoot if i was in math class atm i could help :P

OpenStudy (anonymous):

haha:p

OpenStudy (anonymous):

75 + 2m = 80 + 1.5m 2m - 1.5m = 80 - 75 .5m = 5 m = 10 miles

OpenStudy (skyz):

^ I was thinking along the lives of that

OpenStudy (skyz):

I cant explain it very well doe...

OpenStudy (anonymous):

yes...you just have to set them equal to each other and solve for the miles

OpenStudy (skyz):

Well... That's exactly why your SS is 90 and Mines just 58...